Consider the following abbreviated financial statements for Parrothead Enterprises:

   

PARROTHEAD ENTERPRISES
2017 and 2018 Partial Balance Sheets
Assets Liabilities and Owners’ Equity
2017 2018 2017 2018
  Current assets $ 1,302 $ 1,435 Current liabilities $ 594 $ 637
  Net fixed assets 5,085 6,164 Long-term debt 2,804 2,987

  

PARROTHEAD ENTERPRISES
2018 Income Statement
  Sales $ 16,066
  Costs 7,279
  Depreciation 1,459
  Interest paid 452

  

a. What is owners' equity for 2017 and 2018? (Do not round intermediate calculations.)
b. What is the change in net working capital for 2018? (Do not round intermediate calculations.)
c-1. In 2018, Parrothead Enterprises purchased $2,688 in new fixed assets. How much in fixed assets did Parrothead Enterprises sell? (Do not round intermediate calculations.)
c-2. In 2018, Parrothead Enterprises purchased $2,688 in new fixed assets. What is the cash flow from assets for the year? The tax rate is 24 percent. (Do not round intermediate calculations.)
d-1. During 2018, Parrothead Enterprises raised $584 in new long-term debt. How much long-term debt must Parrothead Enterprises have paid off during the year? (Do not round intermediate calculations.)
d-2. During 2018, Parrothead Enterprises raised $584 in new long-term debt. Wh
Answer #1

As per rules I am answering the first 4 subparts of the question

A: Owners Equity = Total assets- liabilities

2017 = Equity = 1302+5085-594-2804=2989

2018 = Equity = 1435+6164-637-2987 = 3975

B: Change in NWC = Closing NWC- Opening NWC

= (1435-637)-(1302-594)

=$90

C-1: Opening fixed assets+Purchases-sales- depreciation = closing fixed assets

5085+ 2688- Sales-1459 = 6164

Asset Sales = $150

C-2 CF from assets = Operating CF – Capital expenditure- Additions to NWC

= (Sales – cost) – Fixed asset purchases – 90

= (16066-7279) – 2688-90

=6009